### v1

@@ 2.4 Well-posedness
For a.e. ${\boldsymbol{x}}\in\Omega$, since $w({\boldsymbol{x}},0)=w_{0}({\boldsymbol{x}})\leq 1$, we have $w_{\#}({\boldsymbol{x}},0)=0$ and $\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$. If $w({\boldsymbol{x}},t)>e^{-\lambda\,t}$ then $\frac{\partial w_{\#}}{\partial t}=\frac{\partial w}{\partial t}+\lambda\,e^{-\lambda\,t}$, and so
$\frac{\partial w}{\partial t}\,w_{\#}=\Big(\frac{\partial w_{\#}}{\partial t}-\lambda\,e^{-\lambda\,t}\Big)w_{\#},\;\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\;d(w)\,w_{\#}=(\lambda+f)\,e^{-\lambda\,t}\,w_{\#},$ (2.17)
where we used Case 3 of (2.14). On the other hand, if $w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}$ then $w_{\#}({\boldsymbol{x}},t)=0$ and the equalities in (2.17) hold trivially. Putting (2.17) into (2.15) gives for a.e. $t\in I$,
$\int_{\Omega}\Big(\frac{\partial w_{\#}}{\partial t}-\lambda\,e^{-\lambda\,t}\Big)w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}=-\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(\lambda+f)\,e^{-\lambda\,t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}},$
which yields
$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}=\int_{\Omega}\frac{\partial w_{\#}}{\partial t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}=-\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}f\,e^{-\lambda\,t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}\leq 0.$
We conclude that $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}$ is a nonincreasing function of $t$. Thus $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},t)=0$ for a.e. ${\boldsymbol{x}}\in\Omega$, and in turn this yields
$w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }t\in I.$
@@
Suppose that $w_{1}$ and $w_{2}$ are two solutions for PDE II, both bounded between 0 and $e^{-\lambda\,t}$. Then from (2.8) we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
$\int_{\Omega}\frac{\partial w_{1}}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w_{1}\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,w_{1}^{2}+c\,w_{1})\,v\,{\mathrm{d}}{\boldsymbol{x}},$
$\int_{\Omega}\frac{\partial w_{2}}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w_{2}\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,w_{2}^{2}+c\,w_{2})\,v\,{\mathrm{d}}{\boldsymbol{x}}.$
Subtracting the two equations, writing $w_{1}^{2}-w_{2}^{2}=(w_{1}+w_{2})\,(w_{1}-w_{2})$, and taking $v=w_{1}(\cdot,t)-w_{2}(\cdot,t)\in H^{1}(\Omega)$, we obtain
$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|(w_{1}-w_{2})(\cdot,t)\|_{L^{2}(\Omega)}^{2}$
$=-\int_{\Omega}a\,|\nabla(w_{1}-w_{2})|^{2}\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,(w_{1}+w_{2})+c)\,(w_{1}-w_{2})^{2}\,{\mathrm{d}}{\boldsymbol{x}}\leq 0,$
where we used $w_{1}({\boldsymbol{x}},t)\geq 0$ and $w_{2}({\boldsymbol{x}},t)\geq 0$. We conclude that $\|(w_{1}-w_{2})(\cdot,t)\|_{L^{2}(\Omega)}$ is nonincreasing for all $t\in I$, and therefore $\|(w_{1}-w_{2})(\cdot,t)\|_{L^{2}(\Omega)}\leq\|(w_{1}-w_{2})(\cdot,0)\|_{L^{2}(\Omega)}=0$. Hence $w_{1}\equiv w_{2}$ and the solution (bounded between 0 and $e^{-\lambda\,t}$) is unique.
@@ 2.5 A priori bound
For every ${\boldsymbol{y}}\in U$, let $u^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5). For $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}\geq 0$, let $w^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the corresponding solution to PDE II (2.8)–(2.10). Then we have the a priori bounds
$\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)},\qquad C^{\boldsymbol{y}}:=\frac{1+a_{\max}^{\boldsymbol{y}}+\lambda^{\boldsymbol{y}}+f_{\max}}{\sqrt{2\,\min(a_{\min}^{\boldsymbol{y}},\lambda^{\boldsymbol{y}}-\kappa_{\max}^{\boldsymbol{y}})}},$
$\|u^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq\sqrt{\max\big(1+2\,(\lambda^{\boldsymbol{y}})^{2},2\big)}\,e^{\lambda^{\boldsymbol{y}}\,T}\,C^{\boldsymbol{y}}\,\|u_{0}\|_{L^{2}(\Omega)}.$
Recall that the two solutions are related by $u=e^{\lambda\,t}\,w$, with $0\leq u\leq 1$ and $0\leq w\leq e^{-\lambda\,t}$. From PDE II (2.8), we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
$\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,w^{2}+c\,w)\,v\,{\mathrm{d}}{\boldsymbol{x}}.$ (2.18)
Time derivative. Since $0\leq w\leq e^{-\lambda\,t}$, we have $0\leq b\,w^{2}+c\,w\leq(b\,e^{-\lambda\,t}+c)\,w=(\lambda+f)\,w$ and $|(b\,w^{2}+c\,w)v|\leq(\lambda+f)\,w\,|v|$. Thus we obtain from (2.18)
$\Big|\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}\Big|$ $\leq a_{\max}\,\|\nabla w(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\lambda+f_{\max})\,\|w(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
$\leq(a_{\max}+\lambda+f_{\max})\,\|w(\cdot,t)\|_{H^{1}(\Omega)}\,\|v\|_{H^{1}(\Omega)},$
@@
$\|\tfrac{\partial w}{\partial t}(\cdot,t)\|_{H^{1}(\Omega)^{*}}=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}|}{\|v\|_{H^{1}(\Omega)}}\leq(a_{\max}+\lambda+f_{\max})\,\|w(\cdot,t)\|_{H^{1}(\Omega)},$
@@
$\|\tfrac{\partial}{\partial t}w\|_{L^{2}(I;H^{1}(\Omega)^{*})}\leq(a_{\max}+\lambda+f_{\max})\,\|w\|_{L^{2}(I;H^{1}(\Omega))}.$ (2.19)
Gradient. Taking $v=w(\cdot,t)$ in (2.18) and integrating with respect to $t\in I$, we obtain
$\frac{\|w(\cdot,T)\|_{L^{2}(\Omega)}^{2}-\|w(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}=\int_{I}\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t$
$=\int_{I}\int_{\Omega}\frac{\partial w}{\partial t}\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=-\int_{I}\int_{\Omega}a\,|\nabla w|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t-\int_{I}\int_{\Omega}(b\,w^{2}+c\,w)\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t,$
@@
$\frac{\|w(\cdot,T)\|_{L^{2}(\Omega)}^{2}}{2}+\int_{I}\int_{\Omega}a\,|\nabla w|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}(b\,w^{2}+c\,w)\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\frac{\|w(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}.$
Since $0\leq w\leq e^{-\lambda\,t}$, we have $(b\,w^{2}+c\,w)\,w\geq c\,w^{2}$. This yields
$\min(a_{\min},c_{\min})\,\|w\|_{L^{2}(I;H^{1}(\Omega))}^{2}$
$\leq a_{\min}\,\int_{I}\|\nabla w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t+c_{\min}\,\int_{I}\|w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t\leq\frac{\|w_{0}\|_{L^{2}(\Omega)}^{2}}{2},$
with $c_{\min}:=\lambda-\kappa_{\max}>0$, and therefore
$\|w\|_{L^{2}(I;H^{1}(\Omega))}\leq\frac{\|w_{0}\|_{L^{2}(\Omega)}}{\sqrt{2\,\min(a_{\min},\lambda-\kappa_{\max})}}.$ (2.20)
@@
$\|w\|_{{\mathcal{X}}}$ $=\sqrt{\|w\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}w\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}\leq\|w\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}w\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
$\leq(1+a_{\max}+\lambda+f_{\max})\,\|w\|_{L^{2}(I;H^{1}(\Omega))}\leq\frac{1+a_{\max}+\lambda+f_{\max}}{\sqrt{2\,\min(a_{\min},\lambda-\kappa_{\max})}}\,\|w_{0}\|_{L^{2}(\Omega)}.$
Now since $u=e^{\lambda\,t}\,w$, we have
$\|u\|_{{\mathcal{X}}}^{2}$ $=\int_{I}\big(\big\|e^{\lambda\,t}\,w(\cdot,t)\big\|_{L^{2}(\Omega)}^{2}+\big\|\nabla e^{\lambda\,t}\,w(\cdot,t)\big\|_{L^{2}(\Omega)}^{2}+\big\|\tfrac{\partial}{\partial t}[e^{\lambda\,t}\,w(\cdot,t)]\big\|_{H^{1}(\Omega)^{*}}^{2}\big)\,{\mathrm{d}}t,$
where we can take out a factor $e^{2\,\lambda\,t}$ from the first two squared norms on the right-hand side, while the third squared norm can be bounded by
$\big\|\tfrac{\partial}{\partial t}[e^{\lambda\,t}\,w(\cdot,t)]\big\|_{H^{1}(\Omega)^{*}}^{2}$ $=\big\|\lambda\,e^{\lambda\,t}\,w(\cdot,t)+e^{\lambda\,t}\,\tfrac{\partial}{\partial t}w(\cdot,t)\big\|_{H^{1}(\Omega)^{*}}^{2}$
$\leq 2\,e^{2\lambda\,t}\big(\lambda^{2}\,\|w(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}+\|\tfrac{\partial}{\partial t}w(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}\big),$

### v2

@@ 2.4 Well-posedness
For a.e. ${\boldsymbol{x}}\in\Omega$, since $w({\boldsymbol{x}},0)=w_{0}({\boldsymbol{x}})\leq 1$, we have $w_{\#}({\boldsymbol{x}},0)=0$ and $\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$. If $w({\boldsymbol{x}},t)>e^{-\lambda\,t}$ then $\frac{\partial}{\partial t}w_{\#}=\frac{\partial}{\partial t}w+\lambda\,e^{-\lambda\,t}$, and so
$\tfrac{\partial}{\partial t}w=\tfrac{\partial}{\partial t}w_{\#}-\lambda\,e^{-\lambda\,t},\quad\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\quad d(w)\,w_{\#}=(\lambda+f)\,e^{-\lambda\,t}\,w_{\#},$ (2.18)
where we used Case 3 of (2.15). On the other hand, if $w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}$ then $w_{\#}({\boldsymbol{x}},t)=0$ and the equalities in (2.18) hold trivially. Putting (2.18) into (2.16) gives for all $\tau\in I$,
$\int_{0}^{\tau}\!\!\big\langle\tfrac{\partial}{\partial t}w_{\#}-\lambda\,e^{-\lambda\,t},w_{\#}\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\big(a\,|\nabla w_{\#}|^{2}+(\lambda+f)\,e^{-\lambda\,t}\,w_{\#}\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0.$
Using (2.6) then yields
$\frac{\|w_{\#}(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}-\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}=-\int_{0}^{\tau}\!\!\int_{\Omega}\big(a\,|\nabla w_{\#}|^{2}+f\,e^{-\lambda\,t}\,w_{\#}\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq 0.$
Thus $\|w_{\#}(\cdot,\tau)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},\tau)=0$ for a.e. ${\boldsymbol{x}}\in\Omega$, and in turn this yields
$w({\boldsymbol{x}},\tau)\leq e^{-\lambda\,t}\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }\tau\in I.$
@@
Suppose that $w_{2}$ is another solution for PDE II, which is not necessarily bounded between 0 and $e^{-\lambda\,t}$. Subtracting their respective equations analogous to (2.16), we obtain for all $\tau\in I$,
$\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}(w-w_{2}),v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t$
$+\int_{0}^{\tau}\!\int_{\Omega}\big(a\,\nabla(w-w_{2})\cdot\nabla v+b\,(w^{2}-w_{2}^{2})\,v+c\,(w-w_{2})\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0\quad\forall\,v\in{\mathcal{X}}.$
Taking $v=w-w_{2}$ and writing $w^{2}-w_{2}^{2}=2\,w\,(w-w_{2})-(w-w_{2})^{2}=2\,w\,v-v^{2}$, we obtain
$\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}v,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\big(a\,|\nabla v|^{2}+(2\,b\,w+c)\,v^{2}-b\,v^{3}\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0.$
Since $\|v(\cdot,0)\|_{L^{2}(\Omega)}^{2}=0$ and $w\geq 0$, and using (2.6), we conclude that
$\frac{\|v(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}}{2}+a_{\min}\,\|\nabla v\|_{L^{2}([0,\tau];L^{2}(\Omega))}^{2}+c_{\min}\,\|v\|_{L^{2}([0,\tau];L^{2}(\Omega))}^{2}\leq b_{\max}\int_{0}^{\tau}\!\!\int_{\Omega}|v|^{3}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (2.19)
We now prove that $v\equiv 0$ using a proof by contradiction. To facilitate this proof, we introduce the space ${\mathcal{Y}}_{\tau}:=C([0,\tau];L^{2}(\Omega))\cap L^{2}([0,\tau];H^{1}(\Omega))$ for $\tau\in[0,T]$ with the norm
$\|v\|_{{\mathcal{Y}}_{\tau}}^{2}:=\max_{t\in[0,\tau]}\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}+\|\nabla v\|_{L^{2}([0,\tau];L^{2}(\Omega))}^{2}+\|v\|_{L^{2}([0,\tau];L^{2}(\Omega))}^{2}.$
Let $t_{0}:=\max\{\tau\in[0,T]:\|v\|_{{\mathcal{Y}}_{\tau}}=0\}$. Then $v(\cdot,t)=w(\cdot,t)-w_{2}(\cdot,t)=0$ for all $t\leq t_{0}$. We assume by contradiction that $t_{0}<T$. For any $\varepsilon>0$, we take $t_{1}\in(t_{0},T]$ such that $0<\|v\|_{{\mathcal{Y}}_{t_{1}}}<\varepsilon$.
For any $\tau\leq t_{1}$, the right-hand side of (2.19) can be bounded by $b_{\max}\,\theta_{1}^{2}\,\theta_{2}\,\|v\|_{{\mathcal{Y}}_{t_{1}}}^{3}$ using (3.2). Using this to bound each of the three terms on the left-hand side of (2.19) separately and then combining, we deduce that
$\|v\|_{{\mathcal{Y}}_{t_{1}}}^{2}\leq\big(2+\tfrac{1}{a_{\min}}+\tfrac{1}{c_{\min}}\big)\,b_{\max}\,\theta_{1}^{2}\,\theta_{2}\,\|v\|_{{\mathcal{Y}}_{t_{1}}}^{3}<\big(2+\tfrac{1}{a_{\min}}+\tfrac{1}{c_{\min}}\big)\,b_{\max}\,\theta_{1}^{2}\,\theta_{2}\,\,\varepsilon\,\|v\|_{{\mathcal{Y}}_{t_{1}}}^{2},$
@@ 2.5 A priori bound
For every ${\boldsymbol{y}}\in U$, let $u^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5). Then we have the a priori bound
$\|u^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq C^{\boldsymbol{y}}\,\|u_{0}\|_{L^{2}(\Omega)},\quad\mbox{with}\quad C^{\boldsymbol{y}}:=\frac{(1+a_{\max}^{\boldsymbol{y}}+\kappa_{\max}^{\boldsymbol{y}}+f_{\max})\,e^{\kappa_{\max}^{\boldsymbol{y}}T}}{\sqrt{\min(a_{\min}^{\boldsymbol{y}},\kappa_{\max}^{\boldsymbol{y}})}}.$
From PDE I (2.1), we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
$\big\langle\tfrac{\partial}{\partial t}u,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}+\int_{\Omega}\big(a\,\nabla u\cdot\nabla v+(\kappa\,u^{2}+(f-\kappa)\,u)\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}=0.$ (2.20)
Time derivative. Since $0\leq u\leq 1$, we have $|(\kappa\,u^{2}+(f-\kappa)\,u)\,v|=|\kappa\,u\,(u-1)+f\,u|\,|v|\leq(\kappa+f)\,u\,|v|$. Thus we obtain from (2.20)
$\big|\big\langle\tfrac{\partial}{\partial t}u,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\big|$ $\leq a_{\max}\,\|\nabla u(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\kappa_{\max}+f_{\max})\,\|u(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
$\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|u(\cdot,t)\|_{H^{1}(\Omega)}\,\|v\|_{H^{1}(\Omega)},$
@@
$\|\tfrac{\partial u}{\partial t}(\cdot,t)\|_{H^{1}(\Omega)^{*}}=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\langle\tfrac{\partial}{\partial t}u,v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}|}{\|v\|_{H^{1}(\Omega)}}\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|u(\cdot,t)\|_{H^{1}(\Omega)},$
@@
$\|\tfrac{\partial}{\partial t}u\|_{L^{2}(I;H^{1}(\Omega)^{*})}\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|u\|_{L^{2}(I;H^{1}(\Omega))}.$ (2.21)
Gradient. Taking $v=u(\cdot,t)$ in (2.20) and integrating with respect to $t\in[0,\tau]$ for some $\tau\leq T$, and using (2.6), we obtain
$\frac{\|u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}-\|u(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}+\int_{0}^{\tau}\!\!\big(a\,|\nabla u|^{2}+\kappa\,u^{2}+(f-\kappa)\,u\big)\,u\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0,$
@@
$\frac{\|u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}}{2}+\!\int_{0}^{\tau}\!\!\int_{\Omega}\big(a\,|\nabla u|^{2}+f\,u^{2}\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\!\int_{0}^{\tau}\!\!\int_{\Omega}\kappa\,u^{2}\,(1-u)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\frac{\|u_{0}\|_{L^{2}(\Omega)}^{2}}{2}.$ (2.22)
Dropping the nonnegative second term in (2.22) and using $0\leq u\leq 1$, we obtain the inequality
$\|u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}\leq 2\,\kappa_{\max}\int_{0}^{\tau}\|u(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t+\|u_{0}\|_{L^{2}(\Omega)}^{2}.$
The integral form of Grönwall’s inequality (see e.g., [43, Lemma 1.4.1]) then yields
$\|u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}\leq e^{2\,\kappa_{\max}\,\tau}\|u_{0}\|_{L^{2}(\Omega)}^{2},\quad\mbox{and so}\quad\|u\|_{L^{2}(I;L^{2}(\Omega))}^{2}\leq\frac{e^{2\,\kappa_{\max}\,T}-1}{2\,\kappa_{\max}}\|u_{0}\|_{L^{2}(\Omega)}^{2}.$
Keeping only the $a\,|\nabla u|^{2}$ term on the left-hand side of (2.22) and taking $\tau=T$ lead to
$a_{\min}\,\|\nabla u\|_{L^{2}(I;L^{2}(\Omega))}^{2}$ $\leq\kappa_{\max}\,\|u\|_{L^{2}(I;L^{2}(\Omega))}^{2}+\frac{\|u_{0}\|_{L^{2}(\Omega)}^{2}}{2}\leq\frac{e^{2\,\kappa_{\max}\,T}}{2}\,\|u_{0}\|_{L^{2}(\Omega)}^{2}.$
In turn we obtain
$\|u\|_{L^{2}(I;H^{1}(\Omega))}^{2}=\|u\|_{L^{2}(I;L^{2}(\Omega))}^{2}+\|\nabla u\|_{L^{2}(I;L^{2}(\Omega))}^{2}$
$\leq\frac{e^{2\,\kappa_{\max}\,T}-1}{2\,\kappa_{\max}}\|u_{0}\|_{L^{2}(\Omega)}^{2}+\frac{e^{2\,\kappa_{\max}\,T}}{2\,a_{\min}}\,\|u_{0}\|_{L^{2}(\Omega)}^{2}\leq\frac{e^{2\,\kappa_{\max}\,T}}{\min(a_{\min},\kappa_{\max})}\|u_{0}\|_{L^{2}(\Omega)}^{2}.$ (2.23)
@@
$\|u\|_{{\mathcal{X}}}=\sqrt{\|u\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}u\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}\leq\|u\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}u\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
$\leq(1+a_{\max}+\kappa_{\max}+f_{\max})\,\|u\|_{L^{2}(I;H^{1}(\Omega))}\leq\frac{(1+a_{\max}+\kappa_{\max}+f_{\max})\,e^{\kappa_{\max}\,T}}{\sqrt{\min(a_{\min},\kappa_{\max})}}\,\|u_{0}\|_{L^{2}(\Omega)},$